OpenStudy (anonymous):

Find the 11th term in the geometric sequence

OpenStudy (anonymous):

find the 11th term in The geometric sequence in which \[a _{3}\] is 24 and \[a _{5}\] is 96

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

there are two ways to solve this, first method involves finding the common difference and first term

ganeshie8 (ganeshie8):

\(\large a_3 = a r^2 = 24\) \(\large a_5 = a r^4 = 96\)

ganeshie8 (ganeshie8):

can u solve those two equations and find a, r values ?

OpenStudy (anonymous):

no...

ganeshie8 (ganeshie8):

ok let me show u how to solve, take \(\large \frac{a_5}{a_3} \) \(\large \frac{ar^4}{ar^2} = \frac{96}{24}\) \(\large \frac{\cancel{a}r^4}{\cancel{a}r^2} = 4\) \(\large r^2 = 4\) \(\large r = ?\)

ganeshie8 (ganeshie8):

once you have r, find a also then use the nth term formula for geometric sequence

ganeshie8 (ganeshie8):

try...

OpenStudy (anonymous):

r=2?

OpenStudy (anonymous):

@ganeshie8